OpenStudy (anonymous):

the product of a number and 9 increased by 4 is 58 find the number

OpenStudy (anonymous):

9x+4 = 58, x = 6

OpenStudy (anonymous):

Work backwards. 58-4=54 54/9=6 the number is 6

OpenStudy (anonymous):

\[9x+4=58\] \[9x=58-4\] \[9x=54\] \[x=\frac{54}{9} \] \[x=6\]